# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v1

@@ 3.4 Flight tests
Test flights in Evo were conducted on 18th October, 2023. Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation. In total, seven test flights were conducted in the ”Evo-medium” test area and nine test flights in the ”Evo-difficult” test area. In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths. In the ”Evo-medium” test area, the takeoff location was changed after five test flights. In the ”Evo-difficult” test area the takeoff heading was changed after six test flights. The target flight velocity was 1 m/s in all flight tests. The goal was set 36 m forward in the first takeoff location in the ”Evo-medium” test area, 34 m forward in the second takeoff location in the ”Evo-medium” test area, and 42 m forward in the takeoff location in the ”Evo-difficult” test area. The maximum flight height was set to 2.25 m relative to the takeoff location in the ”Evo-medium” test area. For the first test flight in the ”Evo-difficult” test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
@@
Figure 5: Control targets for reference trajectory generation in ”Evo-difficult” forest. Targets were spread widely to the test site to increase the probability that there would be targets visible in the recorded camera data regardless the path that the robot takes.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair were used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, quality setting was set to high meaning that two times downsampled images were used; the settings for number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, ground control points (GCPs) were used in the processing with an accuracy setting of ±$\pm 0.005$ m. Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle-adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
Point clouds were derived using datasets from four flights conducted in the ”Evo-medium” test area, each designated as follows: I: the first flight, II: the second flight, III: the third flight, IV: the fourth flight, and I-IV: combined processing of all four flights (Combined). The processing was carried out using all GCPs. Additionally, to evaluate the impacts of GCPs, point clouds were generated from the first flight trajectory processed with different GCP configurations, denoted as: I0GCP: no GCPs, I1GCP: one GCP at the beginning of the trajectory, I2GCP: two GCPs, one at the beginning and one at the end of the trajectory, and I3GCP: three GCPs, positioned at the beginning, middle, and end of the trajectory.
The experiments were conducted in dense forests, where collecting reference data using GNSS was not possible due to the forest canopy blocking the signal. Reference trajectories were generated for the flights using SfM processing. In each case, the referenece trajectory was derived from the first takeoff heading and orientation in both Evo test environments.
@@ 3.6 Tree parameter estimation
In this study, trees were detected and their DBHs estimated using an automatic stem diameter estimation algorithm originally proposed by Hyyppä et al. (2020b), and further developed by Hyyppä et al. (2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS. The method described by Hyyppä et al. (2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)), and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).
@@ 3.7.1 Autonomous flying performance
The performance and reliability of the autonomous navigation of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
Obstacle avoidance reliability was evaluated based on the mission success. A mission was considered successful if the drone reached the end of the flight without major collisions. Minor touches to thin vegetation or branches were allowed, if the drone was able to continue the mission after that. If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal. In addition to obstacle avoidance, the smoothness of the trajectory was assessed. The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
The accuracy of the VIO localization was assessed by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al. (2012), with respect to the SfM reference trajectories. In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated. ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991). The root mean square error (RMSE) of the individual state errors is then calculated to generate a single metric value representing the entire trajectory’s accuracy.. In the mathematical form the error can be presented as
@@
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the drone’s onboard computer, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
Since the coded targets are easily detectable, unique, and unnatural objects, the loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets. Since there were no coded targets for forming the reference trajectories, only the ability to detect the loops inside dense natural forest was tested, and the accuracy of the loop corrected trajectory was not evaluated. VINS-Fusion estimate with the loop detection algorithm was ran for the collected data with the drone’s onboard computer, and the effect of amount of features in pictures for loop detection capability was examined.
@@ 3.7.2 DBH estimation performance
The accuracy of the DBH estimation from the drone point cloud data was evaluated using Terrestrial Laser Scanning (TLS) data as a reference. The TLS scannings were conducted in April 2020, and tree height, diameter at breast height (DBH), and other parameters were estimated from the dataset. The estimated accuracy of the DBH measurement was less than 1 cm.
Several metrics were used to evaluate the accuracy of the DBH estimation, including Root Mean Square Error (RMSE), Relative RMSE (RMSE%), Bias, Relative Bias (bias%), and Standard Deviation.
@@
Additionally, the completeness of stem detection was evaluated by comparing the total number of reference trees found from the drone-collected point cloud data to the total number of reference trees within the point cloud boundary.
@@
The impact of tree growth between the collection of drone data and reference measurements was not taken into account. Variability in tree growth, influenced by factors such as species, locations, health, age, is acknowledged as a source of error. Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates. Additionally, some reference trees fell between the measurements and were not included in the estimations.
@@ 4.1.1 Obstacle avoidance
The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest. Consistent with these expectations, the trajectories were smoother and the success rate was better in ”Evo-medium” than in ”Evo-difficult” test area. In ”Evo-medium” all performed test flights were successful and five out of seven test flights were performed with smooth trajectories. In ”Evo-difficult” one of nine test flights failed and none of the test flights was performed with a smooth trajectory.
The reason for emergency stops in both environments was late detection of thin and dry spruce branches. In the failed test flight in ”Evo-difficult” the reason was also a collision to a similar thin low branch.
The test flight success rates in both test forests are summarized in the Table 3. Figure 7 presents examples of performed trajectories and obtained point clouds in all test forests.
@@
”Evo-medium” Old Spruce forest 650 34 - 36 7/7 5/7
”Evo-difficult” Young Spruce forest 2000 42 8/9 0/8
(a) ”Evo-medium”
(b) ”Evo-difficult”
Figure 7: Examples of drones own navigation grid maps with performed trajectories (left) and obtained point clouds (right) in (a) ”Evo-medium” and (b) ”Evo-difficult”
@@ 4.1.2 VIO
The trajectory estimate error against the reference trajectory was successfully calculated for five flights in the medium forest and for three flights in the difficult forest. In the difficult forest, one of the flights was performed without coded targets making the reference trajectory non-comparable to others, in one flight there was a gap in the recorded onboard data due to a data recording error during a flight, and in one flight Metashape failed to retrieve the reference path. Average $\text{ATE}_{\text{pos}}$ was 0.50 m in the medium forest and 0.34 m in the difficult forest. The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.06 m in the medium forest and 0.07 m in the difficult forest.
The VINS-Fusion estimates with online $t_{d}$ estimation were calculated afterwards for the same flights. With online $t_{d}$ estimation, average $\text{ATE}_{\text{pos}}$ was 0.47 m in the medium forest and 0.33 m in the difficult forest. The standard deviation between $\text{ATE}_{\text{pos}}$ was 0.02 m in the medium forest and 0.04 m in the difficult forest.
As the small standard deviation between the $\text{ATE}_{\text{pos}}$ estimates indicates, the error between the VINS-Fusion estimates and the reference trajectories was similar in all flights. In every flight, VINS-Fusion was estimating the shape of the trajectory well but underestimating the flight distance. Figure 8 presents an example of VINS-Fusion trajectory estimate and the reference trajectory in the difficult test forest. Table 4 presents the VINS-Fusion estimate errors in all flights.
@@
”Evo-medium” 1 0.5567 0.4816
”Evo-medium” 2 0.5599 0.4790
”Evo-medium” 3 0.4538 0.4769
”Evo-medium” 4 0.4945 0.4722
”Evo-medium” 5 0.4313 0.4400
”Evo-difficult” 1 0.3787 0.3340
”Evo-difficult” 2 0.3853 0.3727
”Evo-difficult” 3 0.2680 0.2962
Figure 9 shows an example of VINS-Fusion tracking image during a test flight in the dense test forest. The algorithm detects stable features to track from all parts of the picture. The image shows also flying dry leaves blown up by the air flow from the propellers of the drone. The moving leaves did not cause visible disturbances to the tracking of the pose.
(a) Top view
(b) Side view
Figure 8: Example of a VINS-Fusion trajectory estimate and a reference trajectory obtained with Agisoft Metashape in the difficult forest from a) top and b) side.
@@
In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests. Figure 10 shows examples of detected loop closure image pairs from both paths. The number of detected loops can be increased by increasing the amount of features per image in the VINS-Fusion configuration. By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path. By decreasing the minimum distance between features to 30 pixels, VINS-Fusion detected a loop closure between 82 image pairs in the shorter path and 63 image pairs in the longer path. However, increasing the number of features also increases the computational burden of the VIO, and the effects on the real-time performance during flying were not tested in this study.

### v2

@@ 3.4 Flight tests
Test flights in Evo were conducted on 18th October 2023. Multiple test flights were performed to evaluate the forest-type effects on the performance of autonomous navigation. In total, seven test flights were conducted in the ’Evo-medium’ test area and nine test flights in the ’Evo-difficult’ test area. In each test environment, either the takeoff location or the takeoff heading was altered during the test flight campaign to force the drone to follow different paths. In the ’Evo-medium’ test area, the takeoff location was changed after five test flights. In the ’Evo-difficult’ test area, the takeoff heading was changed after six test flights. The target flight velocity was 1 m/s in all flight tests. The goal was set 36 m forward in the first takeoff location in the ’Evo-medium’ test area, 34 m forward in the second takeoff location in the ’Evo-medium’ test area, and 42 m forward in the takeoff location in the ’Evo-difficult’ test area. The maximum flight height was set to 2.25 m relative to the takeoff location in the ’Evo-medium’ test area. For the first test flight in the ’Evo-difficult’ test area, the maximum flight height was set to 2.5 m with respect to the takeoff location and was increased to 2.75 m for the rest of the flights.
@@ 3.5 Structure-from-Motion Multi-View Stereo based reconstruction
Stereo image processing was carried out using Agisoft Metashape software (Agisoft LLC, St. Petersburg, Russia) (Metashape Development Team 2021) following similar processing steps as introduced by Viljanen et al. (2018). Every sixth stereo pair was used in the Metashape processing with a stereo baseline of 5 cm and an accuracy setting of ±0.0015 m. In the image orientation processing, the quality setting was set to high meaning that two times downsampled images were used. The settings for the number of key points and tie points per image were 40,000 and 4,000, respectively. To enhance the quality of the post-processed reference trajectories, GCPs were used in the processing with an accuracy setting of ±$\pm 0.005$ m. Automatic outlier removal was performed using gradual selection tools of the software based on re-projection error, reconstruction uncertainty, and projection accuracy. Subsequently, a bundle adjustment was carried out, involving optimization of the sparse point cloud along with the interior (IOP) and the exterior orientation parameters (EOP). Following that, a dense point cloud was calculated using the highest quality setting and mild depth filtering. Finally, the EOPs and dense point clouds were exported.
Point clouds for DBH estimation were derived using datasets from four flights conducted in the ’Evo-medium’ test area, each designated as follows: I: the first flight, II: the second flight, III: the third flight, IV: the fourth flight, and I-IV: combined processing of all four flights (Combined). The processing was carried out using all GCPs. Additionally, to evaluate the impacts of GCPs, point clouds were generated from the first flight trajectory processed with different GCP configurations, denoted as: I0GCP: no GCPs, I1GCP: one GCP at the beginning of the trajectory, I2GCP: two GCPs, one at the beginning and one at the end of the trajectory, and I3GCP: three GCPs, positioned at the beginning, middle, and end of the trajectory.
The experiments were conducted in dense forests, where collecting reference data using GNSS was not possible due to the forest canopy blocking the signal. Reference trajectories were generated for the flights using SfM processing. In each case, the reference trajectory was derived from the first takeoff heading and orientation in both Evo test environments. Figure 6 visualizes the GCPs, stereo images, and reference trees during the point cloud processing in the Agisoft Metashape.
@@ 3.6 Tree parameter estimation
In this study, trees were detected, and their DBHs were estimated using an automatic stem diameter estimation algorithm proposed by Hyyppä et al. (2020b) and further developed by Hyyppä et al. (2020a) to be suitable for a larger variety of point clouds, including photogrammetric point clouds, and point clouds produced with 3D mobile laser scanners and TLS. The method described by Hyyppä et al. (2020a) was applied to estimate the DBH from point clouds collected with the drone and a TLS system, the latter of which provided reference measurements of DBH (see details in Section 3.7).
The point cloud was first segmented into 40-cm-tall (20-cm-tall for TLS) vertical segments, and 5-s-long temporal segments. The temporal segmentation was done to mitigate point cloud distortions caused by potential slowly-accumulating positioning errors of the moving drone system, and thus, the temporal segmentation was not performed for the TLS data. The points potentially corresponding to the tree stems were detected by clustering the points with density-based clustering (DBSCAN, (Ester et al. 1996)) and filtered by fitting a circle to each cluster while accepting only those clusters satisfying the following heuristic criteria: the cluster contained more than 35 points and at least 80 % of the points within 30 mm of the circular arc, the fitted circle radius was between 4 cm and 40 cm, and the points covered at least a 60° central angle of the fitted circle. The accepted clusters were further grouped into tree stems by applying DBSCAN for the center points of the clusters, and the growth direction of each tree was estimated with principal component analysis (PCA). The points within accepted clusters were projected into the plane perpendicular to the growth direction for the final diameter estimation. The diameters were processed into a stem curve by automatically removing clear outliers, followed by fitting a smoothing spline to the remaining diameter estimates as a function of height. Finally, the DBH estimates were calculated by interpolating the stem curve at the height of 1.3 m above the ground. For a detailed description of the method, the reader is referred to Hyyppä et al. (2020b, a).
@@ 3.7.1 Autonomous flying performance
The performance and reliability of the autonomous flight of the prototype were evaluated based on two factors: the reliability of the obstacle avoidance and the accuracy of the VIO localization.
Obstacle avoidance reliability was evaluated based on the mission success. A mission was considered successful if the drone reached the end of the flight without major collisions. Minor touches to thin vegetation or branches were allowed if the drone was able to continue the mission after that. If the given goal was inside an obstacle, the mission was considered successful if the drone managed to fly next to the goal. In addition to obstacle avoidance, the smoothness of the trajectory was assessed. The trajectory was considered smooth if the system did not perform any emergency stops during the flight.
The accuracy of the VIO localization was evaluated by calculating the Absolute Trajectory Error (ATE), as proposed by Sturm et al. (2012), with respect to the post-processed SfM reference trajectories. In this study, ATE was calculated only for the position estimate, and the error in the orientation estimate was not evaluated. ATE was calculated using the trajectory evaluation toolbox by Zhang and Scaramuzza (2018), which aligns the trajectory estimates with the reference trajectories with the method by Umeyama (1991). After aligning the trajectories, the toolbox calculates the RMSE of the individual state errors to generate a single metric error value representing the accuracy of the entire trajectory. In the mathematical form, the error can be presented as
@@
In the first tests conducted in a simulator as described by (Karjalainen et al. 2023), the VINS-Fusion estimate was found to be more accurate when the $t_{d}$ value was fixed to constant, compared to when the online estimation method (Qin and Shen 2018) was used. For this reason, during the test flights, the $t_{d}$ value was fixed to the one obtained from kalibr. However, the calibration estimate is never perfect, and the prototype did not have actual hardware synchronization between the camera and the autopilot IMU. For that reason, after the test flights, the VINS-Fusion estimation was also run with online estimated $t_{d}$ with the camera and IMU data recorded during the flight tests. The estimates were performed with the onboard computer of the drone, and the $t_{d}$ estimate obtained with kalibr was used as the initial value of online estimation.
The loop detection capability of VINS-Fusion was evaluated in a natural forest environment in Paloheinä, without the use of artificial targets that could affect the system performance. Since there were no GCPs for forming the reference trajectories, only the ability to detect the loops inside dense natural forests was tested, and the accuracy of the loop-corrected trajectory was not evaluated. VINS-Fusion estimate with the loop detection algorithm was run for the collected data with the onboard computer of the drone, and the effect of a maximum number of features in pictures for loop detection capability was examined.
@@ 3.7.2 DBH estimation performance
The accuracy of the DBH estimation from the drone point cloud data was evaluated using TLS data as a reference. The TLS scannings were conducted in April 2020, and tree height, DBH, and other parameters were estimated from the dataset. The estimated error of the DBH measurement was less than 1 cm.
Several metrics were used to evaluate the accuracy of the DBH estimation, including RMSE, Relative RMSE (RMSE %), Bias, Relative Bias (bias %), and Standard Deviation.
@@
The completeness of stem detection was evaluated by comparing the total number of reference trees found from the drone-collected point cloud data to the total number of reference trees within the point cloud boundary.
@@
The impact of tree growth between the collection of drone data and reference measurements (four growth seasons) was not taken into account. Variability in tree growth, influenced by factors such as species, locations, health, and age, is acknowledged as a source of error. Considering that the tree growth was relatively constant in the test area, forming a major component of the bias in the estimates. Additionally, some reference trees fell between the measurements and were not included in the estimations.
@@ 4.1.1 Obstacle avoidance
The task of avoiding collisions during the flight is naturally more demanding in a dense forest than in a sparse forest. Consistent with these expectations, the trajectories were smoother and the success rate was better in ’Evo-medium’ than in ’Evo-difficult’ test area. In ’Evo-medium’, all performed test flights were successful and five out of seven test flights were performed with smooth trajectories without any emergency stops during the flight. In ’Evo-difficult’, one of nine test flights failed. None of the test flights in ’Evo-difficult’ was performed with a smooth trajectory. However, despite the emergency stops during the flight, the system was still able to continue the mission after the stops, leading to a successful mission in eight of the nine test flights.
The reason for emergency stops in both environments was the late detection of thin and dry spruce branches. In the failed test flight in ’Evo-difficult’ the reason was also a collision to a similar thin low branch.
The test flight success rates in both test forests are summarized in Table 3. Figure 8 presents examples of performed trajectories and obtained point clouds in both test forests.
@@
’Evo-medium’ Old Spruce forest 650 34 - 36 7/7 5/7
’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/8
(a) ’Evo-medium’
(b) ’Evo-difficult’
Figure 8: Examples of drones’ own navigation grid maps with performed trajectories (left) and obtained point clouds (right) in (a) ’Evo-medium’ and (b) ’Evo-difficult’
@@ 4.1.2 VIO
The trajectory estimate error against the reference trajectory was calculated for the first five flights in ’Evo-medium’ and for three flights in the difficult forest. The flights from the second takeoff location in ’Evo-medium’ and from the second takeoff orientation in ’Evo-difficult’ were not used in the VIO estimation since GCPs used in the reference trajectory processing were positioned along the first flight trajectories in both environments. In ’Evo-difficult’, one of the flights from the first takeoff orientation was performed without GCPs, leading to a non-comparable reference trajectory, one of the flights had a data recording error during a flight, and in one flight, Metashape failed to retrieve the reference path.
Figure 9 shows an example of the VINS-Fusion trajectory estimate compared to the reference in ’Evo-difficult’ while Table 4 summarizes the VINS-Fusion estimation errors across all flights. The average $\text{ATE}_{\text{pos}}$ was 0.50 m in ’Evo-medium’ and 0.34 m in ’Evo-difficult’ with standard deviations of 0.06 m and 0.07 m, respectively, without online $t_{d}$ estimation. The VINS-Fusion estimates with online $t_{d}$ estimation were computed post-flight for the same trajectories, yielding an average $\text{ATE}_{\text{pos}}$ of 0.47 m in ’Evo-medium’ and 0.33 m in ’Evo-difficult’ with standard deviations of 0.02 m and 0.04 m, respectively.
(a) Top view.
(b) Side view.
Figure 9: Example of a VINS-Fusion trajectory estimate and a reference trajectory obtained with Agisoft Metashape in the ’Evo-difficult’ forest from a) top and b) side.
@@
’Evo-medium’ 1 0.5567 0.4816
’Evo-medium’ 2 0.5599 0.4790
’Evo-medium’ 3 0.4538 0.4769
’Evo-medium’ 4 0.4945 0.4722
’Evo-medium’ 5 0.4313 0.4400
’Evo-difficult’ 1 0.3787 0.3340
’Evo-difficult’ 2 0.3853 0.3727
’Evo-difficult’ 3 0.2680 0.2962
The small standard deviations between the $\text{ATE}_{\text{pos}}$ estimates indicated that the error between the VINS-Fusion trajectories and the reference trajectories was similar across all flights. In each case, VINS-Fusion accurately captured the trajectory shape but systematically underestimated the total flight distance.
Figure 10 shows an example of a VINS-Fusion tracking image during a test flight in ’Evo-difficult’. The algorithm detects stable features to track from all parts of the picture. The image also shows flying dry leaves blown up by the airflow from the propellers of the drone. The moving leaves did not cause visible disturbances to the tracking of the pose.
@@
In the loop closure tests, VINS-fusion detected a loop closure between 20 image pairs in the shorter path and four image pairs in the longer path with the same feature tracking parameters (Table 2) that were used in the flight tests. Figure 11 shows examples of detected loop closure image pairs from both paths. The number of detected loops can be increased by increasing the number of features per image in the VINS-Fusion configuration. By reducing the minimum distance between features from 40 to 35 pixels, VINS-Fusion detected a loop closure between 52 image pairs in the shorter path and 23 image pairs in the longer path. By decreasing the minimum distance between features to 30 pixels, VINS-Fusion detected a loop closure between 82 image pairs in the shorter path and 63 image pairs in the longer path. However, increasing the number of features also increases the computational burden of the VIO, and the effects on the real-time performance during flying were not tested in this study.